### v1

@@ 3.2. Including manipulative agents
Thus, the $Q_{i}(t)$ subset of neighbouring agents influencing $i$–th normal agent, such that $Q_{i}(t)\subset\overline{J}_{i}(t)$, can contain both manipulative and normal agents.
@@
We are interested in the study of the opinion dynamics of normal agents $\{x_{i}(t)\}_{i=1}^{N}$ at time $t+1$, which will depend on its own opinion $x_{i}(t)$ at time $t$ and the opinions $z_{j}(t)\in Q_{i}(t)$ of the neighbouring agents, whether normal or manipulative. So, the iterations defined in the previous sections are still valid for calculating the opinion dynamics of the models, taking into account that the set of confidence needs to be extended with all manipulative agents.
@@
We will assume that the manipulators are in a group organized in such a way that all their agents have the same opinion synchronized. With this, they achieve a greater influence on the normal opinions of their confidence interval because their common opinion has a weight of $K$ manipulative agents in the interval.
@@
In addition, we will study the opinion dynamics in simultaneous models because in this way we are avoiding the random component of choosing, for each iteration $t$, a normal agent on which to apply the model. In this case, the model is applied to all normal agents in each iteration.
@@
In simultaneous implementations, the set $Q_{i}(t)$ can be interpreted (specially in the DW–model and Weighted HK models) as the subset of agents having an effective influence on the opinion $x_{i}(t)$ in the interval of time $[t,t+1]$, so some others agents in the confidence interval can interact with the agent $p_{i}$ but they do not have real influence on it.
3.3. The HK–model with manipulators
@@
Consider an agent space $P=\{p_{i}\}_{i=1}^{N}$ with distribution of opinion at time $t$ given by $\{x_{i}(t)\}_{i=1}^{N}$ and $m\in\mathbb{N}$ groups of stubborn or manipulative agents with opinions $\{f_{j}(t)\}_{j=1}^{m}$ at time $t$. We assume that the $j$–th group comprises of $k_{j}\in\mathbb{N}$ agents and the confidence threshold $\varepsilon>0$ is constant for all $i=1,\ldots,N$. Then, the dynamics of this model is ruled by
(3.3) $x_{i}(t+1)=\frac{1}{\#\left(\overline{J}_{i}(t)\right)}\sum_{j\in\overline{J}_{i}(t)}z_{j}(t)=\frac{1}{\#\left(J_{i}(t)\right)+\underset{j>N}{\sum_{j\in\overline{J}_{i}(t)}}k_{j}}\left[\underset{j\leq N}{\sum_{j\in\overline{J}_{i}(t)}}x_{j}(t)+\underset{j>N}{\sum_{j\in\overline{J}_{i}(t)}}k_{j}f_{j}(t)\right]$
@@
In the figures 9 and 10, we illustrate the opinion dynamics of $N=100$ normal agents with an equispaced initial distribution in the opinion space $[-1,1]$ and an only group of 8 manipulators during 200 iterations with a confidence threshold $\varepsilon=0.1$. Each of the six graphics describes the effect of the group of manipulators moving theirs opinions linearly from $f(0)=-0.8$ to $f(t_{\Delta})=0.8$ for $t_{\Delta}\in\{35k\}_{k=1}^{6}$ iterations and, after, $f(t)=0.8$ for all $t>t_{\Delta}$. Observe that the value $t_{\Delta}$ corresponds to the number of iterations in which the manipulators reach their final opinion.
@@
Figure 9. Simultaneous HK–model: Trajectories of an equispaced initial opinion distribution of 100 normal agents and a group of 8 manipulators during 200 iterations with a confidence threshold $\varepsilon=0.1$. The slopes of the increasing segment of the opinion $x=f(t)$ of manipulators correspond to $t_{\Delta}\in\{35k\}_{k=1}^{6}$ iterations to achieve the final opinion $f(t_{\Delta})=0.8$ from the initial $f(0)=-0.8$. The dashed line represents the trajectory of the opinion of the manipulative group.
Figure 9 shows that the manipulative group “pulls” the opinions of normal agents such that, for fixed confidence threshold, numbers of normal and manipulative agents, the number of clusters depends on the speed at which the opinion of manipulators changes. In this model, the group of manipulative agents is not always able to convince the normal agents in its confidence interval because the speed with which it changes its opinion is not always slow enough to attract a large number of normal agents with them, leaving the confidence interval of the group. In the figure 10, the orange bars represent graphics opinion/number of agents of the final clusters at time $T=200$. Observe that, in this example, for manipulators who change their minds at higher speeds, the most populated clusters remain at the opposite side of the final opinion of the group of manipulators and as that speed decreases there is a displacement of the clusters in the same direction as the group does. This effect will be explained in section 3.3.2.
Figure 10. Simultaneous HK–model: Clusters with their number of agents. The black arrow and dashed lines indicate the movement of the opinion of the manipulative agents corresponding to the dynamic of figure 9.
3.3.1. Dynamic with one manipulative group
It is known that, for a symmetric equispaced initial distribution $\mathbf{x}(0)=\{x_{i}(0)\}_{i=1}^{N}$, the mean and the median opinions $\overline{\mathbf{x}}$, $\mathbf{x}_{M}$ in the HK–model (2.3) without any manipulative agent in the opinion space $[-1,1]$ are $\overline{\mathbf{x}}=0$ and $\mathbf{x}_{M}=0$ (see figure 3). The deviation values of the final distribution of opinions $\mathbf{x}(T)=\{x_{i}(T)\}_{i=1}^{N}$ depends on the numbers of final clusters. To analyse the behaviour of the model with a only group of manipulators, we will set the confidence threshold to $\varepsilon=0.1$ to compare the results obtained with a final distribution with several clusters in the case of the dynamics in which only normal agents intervene.
As a first simulation, we will use the initial distribution $\mathbf{x}(0)=\{-1+\frac{2i}{101}\}_{i=1}^{100}$ of normal agents and a group of $K$ manipulators whose opinion changes linearly from $f(0)=-0.8$ to $f(t_{\Delta})=0.8$ and $f(t)=0.8$ for $t\geq t_{\Delta}$. The values of $K$ and $t_{\Delta}$ are chosen to be $K\in\{0,\ldots,30\}$ and $t_{\Delta}\in\{10+6j\}_{j=0,\ldots 30}=\{10,16,22,\ldots,190\}$. The criterion for stopping iterations is
[…]
The aim of this simulation is to observe the behaviour of the mean values of the final distribution of the normal agents as well as its dispersion measured by the standard deviation of the complete final distribution.
In the graph on the left of figure 11, we represent the mean value of the final distribution $\mathbf{x}(200)$ by a color map depending on the number $t_{\Delta}$ of iterations ($x$–axis) and the number $K$ of agents in the manipulative group ($y$–axis). We observe that, for small values of the variables, the effect of the manipulative group is not very significant (in terms of the mean value) since the mean values remain close to the mean achieved without manipulators, i.e. $\overline{\mathbf{x}}\simeq 0$, illustrated in the graph by dark blue colors.
Figure 11. Simultaneous HK–model: Mean value (left), standard deviation (centre) and number of final clusters (right) depending on the number $K$ of manipulators and the number $t_{\Delta}$ of iterations the manipulators are using to reach their final opinion $f(t_{\Delta})=0.8$ from their initial opinion $f(0)=-0.8$. The confidence threshold is fixed at $\varepsilon=0.1$ constant.
Small values of $t_{\Delta}$ imply high speeds in the change of opinion of the manipulators. When the values of $t_{\Delta}$ are increased, the opinion of this group changes more slowly and this allows the manipulators to shift the average value of the opinion of normal agents in the same direction as the opinion of the manipulative group. Bright yellow indicates mean values close to the final opinion of the manipulators.
The graph on the centre of figure 11 shows the standard deviation of the final distribution of normal agents. We can observe that, when the group of manipulators has little influence on normal agents (small values of $K$ and $t_{\Delta}$), the standard deviation of the final distribution takes values $\sigma\simeq 0.55$ slightly higher than the value without manipulators $\sigma=0.5342$ (illustrated in yellow colors). As the variables grow, it can be observed that the standard deviations decrease to $\sigma=0$ (in dark blue), indicating that consensus is reached. This consensus is achieved in the opinions close to the final opinion of the group of manipulators $f(t)=0.8$ for $t\geq t_{\Delta}$. There are no significant differences in the dispersion behaviour if we use the mean absolute deviation, so we only show the standard deviation values $\sigma$.
On the right of figure 11, we can see the number of final clusters, where dark blue represents consensus, that is, all opinions converge to only one cluster.
Figure 12. Simultaneous HK–model: Median value (left), median absolute deviation (centre) and number of iterations in which manipulators behave like stubborn agents after reaching their final opinion $f(t_{\Delta})=0.8$ until the stopping criterion is satisfied (right).
In figure 12, we can observe a similar behaviour of the median value (left) and the median absolute deviation (centre). The graph on the right shows the number of iterations required, after the group of manipulators reaches their final opinion, for the distribution to stabilize.
A complementary way to study how a manipulative group changes the dynamics of opinion is to consider the range of $\alpha$of opinions around the median value, that is, we consider the dynamics of interval $I(t)=[x_{a}(t),x_{b}(t)]$ in the opinion space such that it contains only $\alpha$of the central opinions and excludes $\frac{1}{2}(100-\alpha)$of the most extreme opinions on each side of the opinion space. In our simulation, we set $\alpha=90$% and since $N=100$ normal agents, so once the vector $\mathbf{x}(t)$ is sorted in ascending order, we get the interval $I(t)=[\mathbf{x}_{6}(t),\mathbf{x}_{95}(t)]$. In figure 13, we can observe after the last iteration $t=t_{max}$, the distances of $x_{6}$ and $x_{95}$ to the median value $\mathbf{x}_{M}$ respectively (graphs on the left and center). To study the symmetry of the interval $I$, we compute the difference $d_{-}-d_{+}$ between the distances $d_{-}=|x_{6}-\mathbf{x}_{M}|$ and $d_{+}=|x_{95}-\mathbf{x}_{M}|$. In the graph on the right, the sea green color represents the symmetry of the interval. The blue colors indicate that subinterval $[\mathbf{x}_{M},\mathbf{x}_{95}]$ is larger than the subinterval $[x_{6},\mathbf{x}_{M}]$, while the yellow colors represent the opposite case. It can be seen that, when consensus is not reached, there is a curved blue stripe in which the distance $d_{+}$ is greater than $d_{-}$, which corresponds to a final distribution of two clusters (see figure 13 right). Below that stripe, the three cases are mixed without there being clearly differentiated regions, because the median $\mathbf{x}_{M}$ jumps from one cluster to another with small variations of $t_{\Delta}$ and $K$.
Figure 13. Simultaneous HK–model: Length of the subintervals $[x_{6},\mathbf{x}_{M}]$ (left) and $[\mathbf{x}_{M,}x_{95}]$ (centre) and difference between both length (right).
Thus, it is noted that, when there is intervention of a manipulative group, the central values of the opinion (mean and median) of the normal population move in the same direction as the manipulators (from $\overline{\mathbf{x}}=0$ and $\mathbf{x}_{M}=0$ without manipulators) while the deviations decrease. This tells us that the manipulators manage to shift and narrow Overton’s window on the subject of study. The narrowing in the range of opinions is also due to the nature of the HK–model, since it tends to concentrate opinions on intermediate values.
3.3.2. Analysis of the opinion dynamics among groups of manipulators and normal groups
@@
Notice, as shown in figures 3, 4 and 9, that clusters take shape in each confidence interval after a few iterations, so it is appropriate to analyse the interaction among groups of normal agents and groups of manipulators.
@@
If we consider $m$ manipulative groups with opinion $\{f_{j}(t)\}$ and $K_{j}\in\mathbb{N}$ agents for $j=1,\ldots,m$ interacting with $n$ groups with $N_{s}\in\mathbb{N}$ normal agents with opinion $\{x_{i(s),s}(t)\}$ for $s=1,\ldots,n$ and $i=1,\ldots,N_{s}$ in the same confidence interval then, by iteration (3.3) and for any $\alpha=1,\ldots,n$ and $\beta=1,\ldots,N_{\alpha}$, we have
(3.4) $x_{\beta,\alpha}(t+1)=\frac{\sum_{j=1}^{m}K_{j}f_{j}(t)+\sum_{s=1}^{n}N_{s}x_{i(s),s}(t)}{\sum_{j=1}^{m}K_{j}+\sum_{s=1}^{n}N_{s}}$
@@
and since the left hand side of the expression does not depend on $\alpha$ or $\beta$, then all normal groups reach a local consensus in the first iteration becoming an aggregated normal group of $N=\sum_{s=1}^{n}N_{s}$ agents with the same opinion. Moreover, by the linearity of the mean values, we can simplify the study of the interaction among several groups of normal agents and manipulators to the interaction of just one group of each type. The opinions of such groups will be the mean values of the opinions of all agents in each one. Note that these new groups will be “virtual” because there may not be real agents with its opinion.
@@
So, let us consider a group of $N$ normal agents such that $x_{i}(0)=x_{0}$ for $i=1,\ldots,N$ and a group of $K$ manipulative agents such that their opinions are defined by $f(t)=f_{0}+\lambda t$ with $|f_{0}-x_{0}|<\varepsilon$ and $\lambda\in\mathbb{R}$. Since $x_{i}(0)=x_{j}(0)$ for all $i,j=1,\ldots,N$ then we can do without sub-indexes naming the opinion of normal agents by $x=x_{i}$. Then we have that opinion dynamics follows the formula
@@
A trivial consequence of the proposition 3.1 is the following corollary.
@@
gives us the relationship among the quantities $K$, $N$, $\lambda$ and $\varepsilon$ to ensure that the group of manipulators influences the group of normal agents for all $t>0$. Observe that $\lambda$ describes the speed of change of the opinion of the manipulative group, so when this rate decreases, the ability of the manipulators to influence normal agents increases. Similarly, when the proportion of manipulative agents to the total number of agents (manipulators and normals) is increased, which is given by $w_{f}=\frac{K}{K+N}$, the ability of the manipulator group to influence the normal population is also increased. The same effect has the confidence threshold $\varepsilon$ so, in the HK–model, the more distrustful the population (small values of $\varepsilon$) the lower the influence of the manipulative group. If each normal agent trusts any other opinion (large values of $\varepsilon$), the manipulative group will attract the opinion of the group (or groups) of normal agents to its own, thus achieving the manipulation of the whole society.
@@
Corollary 3.3 implies that, whenever $\left|\lambda\right|>\frac{K\varepsilon}{K+N}$ with $\left|f(t_{0})-x(t_{0})\right|\leq\varepsilon$, there will exist $T\geq 1$ such that $\left|f(t_{0}+T)-x(t_{0}+T)\right|>\varepsilon$. Then, the influence of manipulators stops at $t=t_{0}+T$ leaving the confidence interval of the normal group. This can be seen in figures 9 and 10. The manipulative group attracts normal agents into a cluster which follows the trajectory of opinion of manipulators until the number $N$ of normal agents is big enough not to verify the condition $\left|\lambda\right|\leq\frac{K\varepsilon}{K+N}$. Thus, the manipulative group leaves the confidence interval of the cluster of normal agents whose opinion is not longer modified because there is not another agent in said interval.

### v2

@@ 2.4. A general model with manipulative agents
Thus, the subset $\Omega_{i}(t)\subset\overline{J}_{i}(t)$ of neighbouring agents influencing $i$–th normal agent can contain both manipulative and normal agents.
@@
We are interested in the study of the opinion dynamics of normal agents $\{x_{i}(t)\}_{i=1}^{N}\subset[-1,1]$ at time $t+1$, which will depend on its own opinion $x_{i}(t)$ at time $t$ and the opinions $z_{j}(t)$ of the neighbouring agents for $j\in\Omega_{i}(t)$, whether normal or manipulative.
@@
We will assume that the manipulators are in an organized group such that all their agents have always the same opinion and they change it at every iteration. With this, they achieve a greater influence on the normal opinions of their confidence interval because their common opinion has a weight of $K$ manipulative agents in the interval.
@@
Throughout this article, in addition, we will study the opinion dynamics in simultaneous models because in this way we are avoiding the random component of choosing, for each iteration $t$, a normal agent on which to apply the model. Therefore, each iteration will be applied to all normal agents at the same time.
@@
In simultaneous implementations, the set $\Omega_{i}(t)$ can be interpreted as the subset of agents having an effective influence on the opinion $x_{i}(t)$ in the interval of time $[t,t+1]$, so some others agents in the confidence interval can interact with the agent $p_{i}$ but they do not have real influence on it.
Thus, we can study the influence of a group of manipulators on the opinion of normal agents according to the dynamics of diverse variants of the classic models such as Hegselmann–Krause [9], Deffuant–Weisbuch [5] as well as the models introduced by Toccaceli et al. [18].
For this purpose, we will use two different metrics to assess the shift in the opinions of normal agents: the mean/standard deviation of final opinion and the interval defined by primary clusters. The former takes into account all the opinions of the population, although it does not adequately distinguish the clustering.
The latter considers only the range of the largest clusters, allowing the elimination of isolated opinions or small clusters at the extremes of the range of opinions. This second metric [8] consists in assigning an effective weight [12] to each cluster of the final opinion distribution and discarding those which weight is smaller than a screening value $\delta>0$. The resulting interval containing non–discarded clusters defines a central point and a width, that we will call center and amplitude respectively, which we use to compute the influence of manipulators.
Following [8], in order to calculate the effective weights, we will smooth the final distribution by Gaussian kernel smoothers
$G_{k}(x,r_{k})=e^{-\left(\frac{x-r_{k}}{\alpha\varepsilon}\right)^{2}}$
where $\varepsilon$ will be the confidence threshold and $r_{k}=-1+\frac{2k-1}{h}$ are the middle points of the intervals $[\frac{-h+2(k-1)}{h},\frac{-h+2k}{h}]$ for $k=1,\ldots,h$. Then, the value of the smooth distribution $S$ at $r_{k}$ is
$S(r_{k})=\sum_{j=1}^{N}G_{k}(r_{k},x_{j}).$
Once we have determined the indexes $\{k_{1},\ldots,k_{s}\}$ such that $S(r_{k_{i}})$ are local maximums of $S$ for $i=1,\ldots,s$, then the weight of the cluster at $r_{k_{i}}$ is
$W_{k_{i}}=\frac{w_{k_{i}}}{\sum_{i=1}^{s}w^{2}_{k_{i}}}$
where $w_{k_{i}}=\frac{S(r_{k_{i}})}{\sum_{i=1}^{s}S(r_{k_{i}})}$. Finally, we will discard the clusters $r_{k_{i}}$ such that $W_{k_{i}}<\delta$. The remaining clusters, called primary clusters and denoted by $\{\overline{r}_{1},\ldots,\overline{r}_{b}\}$, define an interval $Q=[\mathrm{min}\{\overline{r}_{j}\},\mathrm{max}\{\overline{r}_{j}\}]\subset[-1,1]$ such that its center $\mathbf{c}_{b}$ and its amplitude $\mathbf{a}_{b}$ given by
$\mathbf{c}_{b}=\frac{1}{2}(\mathrm{max}\{\overline{r}_{j}\}+\mathrm{min}\{\overline{r}_{j}\})$\\ $\mathbf{a}_{b}=\frac{1}{2}(\mathrm{max}\{\overline{r}_{j}\}-\mathrm{min}\{\overline{r}_{j}\})$\end{tabular}\right.$
allow to measure the influence of manipulators. We will implement this method with values $h=200$ and $\alpha=0.1$.
There are two features of the manipulative agents that we will vary in this article: the number of manipulators $K$ and the speed at which manipulators change their opinion to influence normal agents. One expects that, for a group of manipulators, the greater their number or the slower their speed, the greater the influence they can have on the other agents.
Notice that throughout this article, we will consider the confidence thresholds of all agents to be equal, that is $\varepsilon_{i}=\varepsilon$ for all $i=1,\ldots,N$.
3. The Hegselmann–Krause model
The first case we will study is the Hegselmann–Krause model [9] or HK–model for short. For the agent $p_{i}$, we consider the set $\Omega_{i}$ to be the whole confidence set $J_{i}$ and the new opinion $x_{i}$ of the agent $p_{i}$ is just the arithmetic mean values of all agents in $J_{i}$. So, we get
(3.1) $x_{i}(t+1)=\frac{1}{|J_{i}(t)|}\sum_{j\in J_{i}(t)}x_{j}(t)$
Figure 1. Simultaneous HK–model: opinion dynamics for an initial distribution of 100 equispaced points $x_{i}(0)=-1+\frac{2i}{101}$ for $i=1,\ldots,100$. The constant values of $\varepsilon$ are specified above every graph.
To illustrate the opinion dynamics in this model, in figure 1, we consider three simulations for an equispaced initial distribution $\{x_{i}(0)\}_{i=1}^{N}\subset[-1,1]$ for $N=100$ agents, with different constant confidence thresholds $\varepsilon\in\{0.1,0.45,0.6\}$ and $T=20$ iterations. Observe that, for small values of $\varepsilon$, clusters appear at the final distribution $\{x_{i}(T)\}_{i=1}^{N}$ whereas for larger values of $\varepsilon$, the number of cluster decreases to 1, that is, consensus is reached. Note also that for intermediate values of $\varepsilon$, some clusters may appear that are not stable over time and end up grouping, reaching a consensus, after a relatively large number of iterations (case $\varepsilon=0.45$ in the center of figure 1), see [9].
Figure 2. Simultaneous HK–model: opinion dynamics for the same initial distribution as figure 1 with one group of 20 stubborn agents each with opinion at $x=0.75$ (marked with a dashed line).
3.1. The HK–model with manipulative agents
@@
Consider an agent space $P=\{p_{i}\}_{i=1}^{N}$ with distribution of opinion at time $t$ given by $\{x_{i}(t)\}_{i=1}^{N}$ and $K$ manipulative agents with opinions $\{f_{j}(t)\}_{j=1}^{K}$ at time $t$. Then, the dynamics of this model is ruled by
(3.2) $x_{i}(t+1)=\frac{1}{|\overline{J}_{i}(t)|}\sum_{k\in\overline{J}_{i}(t)}z_{k}(t)=\frac{1}{|J_{i}(t)|+\underset{k>N}{\sum_{k\in\overline{J}_{i}(t)}}1}\left[\underset{k\leq N}{\sum_{k\in\overline{J}_{i}(t)}}x_{k}(t)+\underset{k>N}{\sum_{k\in\overline{J}_{i}(t)}}f_{k}(t)\right]$
@@
As mentioned in section 2.3 above, we can use the HK–model to illustrate the shift of the Overton window. In figure 3, we show three simulations in the opinion space $[-1,1]$ in which we consider the interval $[-0.7,0.7]$ as the Overton window. We choose an equispaced initial distribution of $N=100$ normal agents with acceptable opinions within the interval $[-0.6,0.6]$ and fix the confidence threshold to $\varepsilon=0.1$. The yellow area represents the Overton window along the time $t>0$. The graphic on the left describes the evolution of opinions without manipulative agents. In the other graphics, one group of $K=10$ (center) and $K=15$ (right) manipulators respectively, intervenes in the iteration. The opinion of all manipulative agents moves linearly from $f(0)=-0.6$ to $f(80)=1$. Note that in the second and third cases, the range of opinions shifts upwards to $[-0.06,0.62]$ and $[0.13,1]$ respectively. Therefore, the Overton windows must also have shifted in the same direction. The graph on the right shows that the manipulators are able to push the opinion of 21 normal agents to the extreme $x=1$ of the opinion space, outside the initial Overton window.
@@
Figure 3. Simultaneous HK–model: dynamics of an equispaced initial opinion distribution of 100 normal agents in $[-0.6,0.6]$ with a confidence threshold $\varepsilon=0.1$. The pictures represent the opinion evolution without manipulative agents (left), with a group of $K=10$ manipulators (center) and a group of $K=15$ manipulator (right). The manipulative group changes its opinion linearly from $f(0)=-0.6$ to $f(80)=1$. The yellow area represents the initial Overton window and the dashed line the trajectory of the opinion of the manipulative group.
3.1.1. Analysis of the opinion dynamics between manipulative and normal groups
The effect of manipulators on normal agents in the HK–model can be explained by the analytical relationship between the parameters involved and this relationship also helps understanding in similar models.
@@ 3.1.1. Analysis of the opinion dynamics between manipulative and normal groups
Notice, as shown in figures 1, 2 and 3, that clusters take shape in each confidence interval after a few iterations, so it is appropriate to analyse the interaction among groups of normal agents and groups of manipulators.
@@
Let us consider $m$ manipulative groups with opinion $\{f_{j}(t)\}$ with $K_{j}\in\mathbb{N}$ agents each for $j=1,\ldots,m$ interacting with $n$ groups with $N_{i}\in\mathbb{N}$ normal agents with opinion $\{x_{i}(t)\}$ for $i=1,\ldots,n$ in the same confidence interval then, by iteration (3.2) of simultaneous HK–model, for any $\alpha=1,\ldots,n$ we have that the opinion of an agent of the normal group $i=\alpha$ is modified by
(3.3) $x_{\alpha}(t+1)=\frac{\sum_{j=1}^{m}K_{j}f_{j}(t)+\sum_{i=1}^{n}N_{i}x_{i}(t)}{\sum_{j=1}^{m}K_{j}+\sum_{i=1}^{n}N_{i}}.$
@@
Since, in simultaneous way, every agent in the confidence interval modifies its opinion under the iteration, the right hand side of (3.3) is the same for all $\alpha$, so it does not depend on $\alpha$, then all normal groups reach a local consensus in the first iteration becoming an aggregated normal group of $N=\sum_{i=1}^{n}N_{i}$ agents with the same opinion. Moreover, by the linearity of the mean values, we can simplify the study of the interaction among several groups of normal agents and manipulators to the interaction of just one group of each type. The opinions of such groups will be the mean values of the opinions of all agents in each one. Note that these new groups will be virtual because there may not be real agents with its opinion.
@@
So, let us consider one group of $N$ normal agents such that $x_{i}(0)=x_{0}$ for $i=1,\ldots,N$ and one group of $K$ manipulative agents such that their opinions are defined by $f(t)=f_{0}+\lambda t$ with $|f_{0}-x_{0}|<\varepsilon$ and $\lambda\in\mathbb{R}$. Since $x_{i}(0)=x_{j}(0)$ for all $i,j=1,\ldots,N$ then we can do without sub-indexes naming the opinion of normal agents by $x=x_{i}$. Then we have that opinion dynamics follows the formula
@@
A trivial consequence of propositions 3.1 and 3.2 is the following corollary.
@@
gives us the relationship among the quantities $K$, $N$, $\lambda$ and $\varepsilon$ to ensure that the group of manipulators influences the group of normal agents for all $t>0$. Observe that $\lambda$ describes the speed of change of the opinion of the manipulative group, so when this rate decreases, the ability of the manipulators to influence normal agents increases. Similarly, when the proportion of manipulative agents to the total number of agents (manipulators and normals) is increased, which is given by $w_{f}=\frac{K}{K+N}$, the ability of the manipulative group to influence the normal population is also increased. The same effect has the confidence threshold $\varepsilon$ so, in the HK–model, the more distrustful the population (small values of $\varepsilon$) the lower the influence of the manipulative group. If each normal agent trusts any other opinion (large values of $\varepsilon$), the manipulative group will attract the opinion of the group (or groups) of normal agents to its own, thus achieving the manipulation of the whole society.
@@
Corollary 3.3 implies that, whenever $\left|\lambda\right|>\frac{K\varepsilon}{K+N}$ with $\left|f(t_{0})-x(t_{0})\right|\leq\varepsilon$, there will exist $T\geq 1$ such that $\left|f(t_{0}+T)-x(t_{0}+T)\right|>\varepsilon$. Then, the influence of manipulators stops at $t=t_{0}+T$ leaving the confidence interval of the normal group. This feature is shown in figures 4 and 5, in which the manipulative group moves its opinion from $f(0)=-1$ to $f(t_{\Delta})=0.75$ in $t_{\Delta}\in\{70,140,210\}$ iterations before it becomes stubborn. So, $\lambda\in\{0.025,0.0125,0.0083\}$ and the manipulators attract normal agents into a cluster which follows the trajectory of opinion of manipulators until the number $N$ of normal agents is big enough not to verify the condition $\left|\lambda\right|\leq\frac{K\varepsilon}{K+N}$. Thus, the manipulative group leaves the confidence interval of the cluster of normal agents whose opinion is not longer modified because there is not another agent in said interval.
Figure 4. Simultaneous HK–model: initial opinion distribution of 100 normal agents in $[-1,1]$ with a confidence threshold $\varepsilon=0.1$ and a group of 8 manipulators changing their opinion linearly from $f(0)=-1$ to $f(t_{\Delta})=0.75$ for $t_{\Delta}\in\{70,140,210\}$ iterations keeping it constant for $t\geq t_{\Delta}$. The dashed line is the trajectory of the opinion of the manipulative group.
Figure 5. Simultaneous HK–model: size of the cluster in the final distribution of the same simulation of figure 4. The manipulators groups the opinions of neighbouring normal agents along its way until the cluster is big enough not to satisfy the condition (3.5).
3.1.2. Simulations for HK–model with one manipulative group
In order to study the behaviour of the model with an only group of manipulators, we will compare the central points and widths of the final distributions using the metrics of the mean/standard deviation and the interval of primary clusters as a function of the number $K$ of agents in the manipulative group and the speed $\lambda$ at which they change their opinion.
Note that, since the ratio of the variation $\lambda$ of the manipulators’ opinion and the number of iterations $t_{\Delta}$ in which the manipulative group change its opinion are related by $\lambda=\frac{f(t_{\Delta})-f(0)}{t_{\Delta}}$, we will use the variable $t_{\Delta}$ instead of $\lambda$.
Fixed the confidence threshold to $\varepsilon=0.1$, we consider an equispaced initial distribution $\mathbf{x}(0)=\{-1+\frac{2i}{101}\}_{i=1}^{100}$ of normal agents and a group of $K$ manipulators whose opinion changes linearly from $f(0)=-1$ to $f(t_{\Delta})=1$ and then $f(t)=1$ for $t\geq t_{\Delta}$. The values of $K$ and $t_{\Delta}$ are chosen to be $K=0,\ldots,30$ and $t_{\Delta}\in\{10j\}_{j=0,\ldots 30}$. The criterion for stopping iterations is
[…]
In figure 6, we represent the mean value, standard deviation and number of clusters (maybe isolated opinions) of the final distribution of normal opinions $\mathbf{x}$ by a color map depending on the number of iterations $t_{\Delta}$ and the number $K$ of agents in the manipulative group. We observe that, for small values of the variables, the effect of the manipulative group is not very significant and the three values remain close to the achieved without manipulators.
Figure 6. Simultaneous HK–model: mean value (left), standard deviation (center) and number of final clusters (right) depending on the number $K$ of manipulators and the number $t_{\Delta}$ of iterations the manipulators are using to reach their final opinion $f(t_{\Delta})=1$ from their initial opinion $f(0)=-1$. The confidence threshold is fixed at $\varepsilon=0.1$ constant.
Small values of $t_{\Delta}$ imply high speeds in the change of opinion of the manipulators. When the values of $t_{\Delta}$ are increased, the opinion of this group changes more slowly and this allows the manipulators to shift the opinions of normal agents in the same direction as the opinion of the manipulative group. In these cases the mean values are close to the final opinion of the manipulators (bright yellow in the left graph) meanwhile the standard deviations decrease to $\sigma=0$ (dark blue in the center graph), indicating that consensus is reached.
Figure 7. Simultaneous HK–model: features of the interval of primary clusters. Center (left), amplitude (center) and number of primary clusters obtained in the same simulation as in figure 6.
In figure 7, we can observe the results of the same simulation when we consider the center, amplitude and number of primary clusters for a screening value $\delta=0.5$ (see figure 8 for frequency of effective weights). A similar behaviour is obtained but the areas of intermediate values of the center and amplitude are slightly wider than for the mean and standard deviation values at the figure 6.
Figure 8. Simultaneous HK–model: histogram of effective weights in simulation of figures 6 and 7.
Thus, it is noted that, when there is intervention of a manipulative group, the central values of the opinion (mean and center) of the normal population move in the same direction as the manipulators (from $\overline{\mathbf{x}}=0$ without manipulators) while the standard deviations and amplitude decrease. This tells us that the manipulators manage to shift and narrow Overton’s window on the subject of study. The narrowing in the range of opinions is also due to the nature of the HK–model, since it tends to concentrate opinions on intermediate values.
4. The Deffuant–Weisbuch model
The following model we will study is the Deffuant–Weisbuch model [5], or DW–model for short.
In this model, for each agent $p_{i}$ we choose the set $\Omega_{i}=\{i,j\}$ where $j\in\{1,\ldots,n\}$ is a random index such that $j\neq i$, then we consider the iteration
Again, we will implement this model in simultaneous way, i.e., each agent performs the iteration (4.1) for each $t$.
In figure 9, examples of opinion dynamics for a simultaneous DW–model are shown. The distribution of opinions is considered stable whenever the values in each cluster differ less than $5\cdot 10^{-4}$ units achieved in $t_{max}$ iterations.
Figure 9. Simultaneous DW–model: opinion dynamics for an initial distribution of $N=100$ equispaced points $x_{i}(0)=-1+\frac{2i}{101}\in[-1,1]$ with $\varepsilon_{i}=\varepsilon\in\{0.1,0.45,0.6\}$ constant for all $i=1,\ldots,100$.
If a distribution of opinion interacts with a group of stubborn agents with constant opinion $x(t)=c\in[-1,1]$ for all $t\geq 0$, then the opinion of agents in the confidence interval of the stubborn agents converges to this group’s opinion. A simulation of this situation can be seen in figure 10 for the same initial distribution as in figure 9 and a group of 20 stubborn agents (marked by a dashed line) with constant opinion $x(t)=0.75$ for all $t\geq 0$. Observe that the stubborn group attracts the neighbouring normal agents to its opinion.
Figure 10. Simultaneous DW–model: opinion dynamics for the same parameters as in figure 9 with a group of 20 stubborn agents with opinion $x=0.75$. The opinion of the stubborn group is shown as a dashed line.